# Studying Workarounds in Software Forms: An Experimental Protocol

### v1

@@ 3.0.3 Metrics
Also, systematicity score indicates how consistently participants apply the same workaround type for the same misalignment type, calculated as the proportion of instances where the most common workaround was used, averaged across misalignment types.
4 Pilot Experiment
We conducted a sample pilot experiment [26] to instantiate our proposed method for studying user workarounds in software applications by analyzing interactions with forms for entering misaligned data. We used convenience sampling, sending invitations to research groups and communities within our network. No prior knowledge of any specific software was required, and 15 participants completed the experiment. They were tasked with entering prepared datasets into instrumented software applications we developed. Across the 3 tasks, each participant encountered at least 4 misalignment types for a total 45 misaligned data.
This pilot experiment allowed us to assess both the feasibility of our experimental method and the concepts we proposed for investigating workarounds. While our findings are not generalizable due to the small sample size, they provided valuable insights.
We observed when users encountering misalignments resorted to workarounds, how these workarounds affected data quality, and how users prioritized their strategies based on the quality of the final data they could enter. We also gained experience with our widget categorization: Open widgets provided flexibility, enabling users to enter unstructured or misaligned data; Predetermined widgets were more prone to data loss due to their rigidity; and how Controlled widgets influenced workaround strategies and data quality. Additionally, we explored whether users developed systematic strategies after repeatedly encountering the same type of misalignment.
Our results indicate that applying this experimental method, along with our proposed research concepts, metrics, and recommendations on instruments and process, can yield valuable insights into workarounds. Conducting larger experiments with more participants, case studies, or natural experiments based on our experimental method would be a worthwhile next step.
4.1 Threats to Validity
Our workaround quality measurement, though based on data retention to reduce subjectivity, may not fully reflect true quality. The systematicity score assumes repeated use of the same workaround type indicates intentional behaviour, potentially overlooking factors like user background or problem interpretation. Participants might attempt more workarounds than usual due to the Hawthorne effect [1]. To mitigate this, we carefully worded experiment instructions to focus on data entry and software interaction without mentioning workarounds.
Despite task and misalignment randomization, earlier tasks may unintentionally improve performance in later ones. The tasks and application domains in our pilot experiment cannot fully represent real-world scenarios or capture the complexity of real-world misalignments, limiting result generalizability. While forms provide a suitable focus for this research stage given their pervasiveness in software, this scope may limit findings’ applicability to systems where workarounds differ.
Our participants’ engineering and computer science backgrounds may have influenced their approach despite being assigned non-technical tasks. For instance, computational thinking [25] might affect workaround types employed. Future experiments should include more diverse participants. Researchers adopting our experimental method should consider participant background diversity.
Finally, our interpretations of participants’ workaround choices may involve subjective judgment. We mitigated this by interpreting results and then cross-referencing participants’ exit survey responses to corroborate findings.
5 Summary and Discussion
We proposed an experimental method to study how users employ workarounds when software forms do not align with their data or goals. A pilot experiment validated its feasibility and research questions. Applying this method provides insights that support future research and help designers better address user needs. Future work could extend this approach to non-form-based systems and explore automated workaround detection to improve user-centred software.
Experiments based on this method can reveal patterns in how users handle form data misalignments. Researchers can determine whether users choose workarounds over abandonment, prioritize data integrity in their workaround strategies, or combine multiple workarounds. Another key aspect is whether users become more efficient after repeatedly encountering misalignments. Additionally, our widget categorization provides a framework for examining the relationship between workaround quality and widget affordances.
Workarounds highlight design flaws but also pose risks to data integrity. Software designers must proactively understand workarounds to identify unmet needs and mitigate risks. Survey results from these experiments can help researchers assess whether detecting workarounds uncovers design problems or confirms that misaligned designs push users toward shadow systems or complete bypass [9].

### v2

@@ 3.0.5 Metrics
Systematicity score indicates to what extent participants consistently applied the same workaround type when encountering the same misalignment type within the same app. For each unique misalignment type, we calculate the proportion of instances where the most common workaround type was used. Since each participant encountered three types of misalignment, with five instances of each per app, we averaged the scores across the three misalignment types to compute a single systematicity score for each participant per app.
4 Results
4.1 Characteristics and frequency of workarounds (RQ1)
Figure 2: Characteristics of workaround types
Our analysis discovered cases where participants used multiple workaround types at once to address a single misalignment within the same record. In these cases, we treated each workaround type separately, totaling of $920$ workarounds. Fig. 2 presents the characteristics of workaround types for all these records.
As shown in Fig. 2(a), participants almost always attempted workarounds when encountering a misalignment. Five participants (22 records) ignored Outlier, Ambiguous, Excess, and Nonstandard misalignments while using REG and open and controlled TA. The most common workarounds are Functional and Data Adjustments, meaning that participants adjusted application functionalities (e.g., establishing conventions or redefining widgets meaning) or manipulated the provided data to align with application constraints. In some cases, they employed Fallback methods, such as using a comment text box (e.g., global feedback) to describe the issue. Data Loss strategies, like truncation, were typically used as a last resort when no other workaround was feasible.
Fig. 2(b) shows instances where multiple workaround types were combined. For the 245 records with more than one workaround type, the most frequent combination was Data Adjustment and Fallback. This suggests that participants found the need to clarify the intent behind Data Adjustments to address subjective issues. In some cases, participants combined Data and Functional Adjustments or multiple Functional Adjustments to achieve accurate data entry. For example, in TA, participants repurposed the student feedback box and checkbox to tag records as suspicious, despite their original purpose. For simpler scenarios, they truncated data for convenience or due to application constraints (e.g., entering only the first out of three phone numbers when the widget allowed for only one). These limitations were occasionally documented using Fallback methods. Fallback and Functional Adjustment are more versatile, frequently co-occurring with other workaround types.
4.1.1 Workarounds and misalignments
We conducted a Chi-Squared test to evaluate the association between workaround types and misalignment types. The results indicated a statistically significant association $(\chi^{2}=295.11,df=20,p<0.000)$. To measure the strength of this association, we calculated Cramer’s V, which revealed a moderate effect size $(V=0.283)$. We conducted post hoc comparisons using pairwise Chi-Squared tests with Bonferroni correction $(p<0.008)$. Fig. 2(c) presents the distribution of workaround types across misalignment types. Significant associations are labelled in bold. In the following, we describe what workarounds participants employed for each misalignment type.
Excess: participants primarily relied on Functional Adjustments to adapt widget functionalities for handling data beyond the app’s native capabilities. Data Adjustments, though less frequent, were used to manipulate data to fit within the app’s existing widgets. Explicit: participants often resorted to Fallback methods to explain the Functional and Data Adjustments they applied to address the misalignment. Nonstandard: participants primarily adjusted the data to normalize it for the app. Some attempts led to data loss, while others are complemented by a Fallback method to explain their approach. Outlier: participants generally followed similar patterns as with Nonstandard misalignments. However, they sometimes chose not to attempt any workarounds. This suggests that, while participants may tolerate minor deviations from standard data, they are more likely to ignore data they perceive as nonsensical. Implicit: participants primarily adjusted the data to perform minor modifications, such as correcting a date format, even if the data was already aligned with the app. In some cases, they entered the data as-is but adjusted the functionality of a nearby widget to report the Implicit data, like repurposing the student feedback text box to report bonus points in TA. Ambiguous: participants interpreted these misalignments in various ways, often choosing to either manipulate the data or adjust widget functionality within the app. This often led to data loss, as they retained only what they considered essential, such as keeping only one phone number in HR while discarding others they deemed unnecessary.
4.1.2 Quality of workarounds
[…]
To validate this hypothesis, we conducted a Spearman’s rank-order correlation test for the relationship between workaround type (treated as ordinal) and workaround quality. The results showed a strong, significant positive correlation $(r_{s}=0.757,p<0.000,n=898)$. This suggests that as participants transition from non-strategic data manipulation or simple Fallback methods (e.g., descriptive comments) to employing data or Functional Adjustments, the quality of input data improves. Based on this and what we presented in Fig. 2(b), Fallback methods are most effective when used alongside Functional or Data Adjustments. When used alone, they tend to produce lower-quality data.
Answer to RQ1: When they encounter misalignments, users almost always employ a workaround. While workarounds result in varying levels of data quality, users generally choose approaches that maximize data integrity. They prioritize Functional and Data Adjustments, as these methods typically yield higher-quality outcomes. In some cases, users opt for simpler alternatives, such as reporting issues in comment boxes or truncating data.
4.2 Widget usage in workarounds (RQ2)
Figure 3: Widgets in workarounds ($n=898$)
As described in Section 3, there are three categories of app widgets relative to how much user input freedom they offer: Open, Predetermined, and Controlled. Across the $898$ records with workarounds (excluding the NW), participants relied on text boxes for the Open widgets. They attempted workarounds with radio buttons, combo boxes, and checkboxes for the Predetermined category. Finally, they used number spinners, sliders, and date pickers for the Controlled category.
4.2.1 Widgets involved in workarounds
Fig. 3(a) presents a heat map of widget usage across different types of workarounds. Text boxes are most frequently used for Fallback workarounds (19.3%), as participants often rely on Open widgets to input misaligned data in an unstructured manner. They also complement other workaround types, such as Functional Adjustment or Data Adjustment, by adding details or clarifying intentions. Occasionally, participants used text boxes for unintended purposes (i.e., Functional Adjustment). Sometimes, we observed Data Loss with Open text boxes, often stemming from participants’ misinterpretation of the text box’s intended purpose. For example, although HR provides a single text box capable of accommodating multiple phone numbers, participants tend to input only one number. This behaviour occurs more frequently than attempting to adjust the data (e.g., entering multiple numbers as a comma-separated value).
Among the Predetermined widgets used for workarounds, participants predominantly employed checkboxes for Functional Adjustments (14.4%). They used various combinations of checkboxes to represent specific concepts or data that could not be entered directly, or repurposed a checkbox to convey different meanings. For example, in HR, some participants selected “fan” as an item under employee amenities, even though the provided data had specified “A/C”. Unlike dropdown menus, combo boxes are editable and allowed some participants to store multiple values, such as multiple topics of interest in REG. Others assumed combo boxes accepted only one value, selecting a single option and causing Data Loss. Radio buttons are less suitable for Functional Adjustments and lead to higher Data Loss, while checkboxes offer more flexibility by enabling multiple selections to approximate intended input. Data Adjustment with radio buttons and Fallback use of checkboxes were rare (0.6%), with the former involving selecting the closest option (e.g., ”Pasta (veg)” for ”Pasta with meat” in REG) and the latter activating a text box for comments.
Data Adjustments were more common with Controlled widgets, especially date pickers, sliders, and number spinners. While these widgets enforce constraints, the results suggest that they they provide more flexibility than Predetermined widgets. Participants creatively interpreted constraints, modified input formats, or manipulated data to work within widget restrictions while entering needed data. These adjustments were often complemented by Fallback entries in Open widgets.
4.2.2 Impact of widgets workaround quality
Fig. 3(b) shows the quality of attempted workarounds across widget categories. We observed more frequently higher-quality workarounds (3–4) across all widget categories than lower-quality (1–2). This suggests that participants preserved data integrity despite widget restrictions by employing creative solutions, such as repurposing fields to capture additional information. For instance, they modified a combo box value to include job seniority in HR. entered a bonus grade in an irrelevant text box in TA, and used the “allergic” checkbox, even though not marked in the data, with the meal preference checkbox in REG to indicate dietary preferences.
We observe a higher rate of data quality scores of 3 in Controlled widgets. This is because users often adhere to the constraints of these widgets and, when necessary, use complementary methods to capture data that cannot be directly entered due to the widget’s controls. In contrast, Predetermined widgets show a higher rate of lower-quality data with a score of 2. This is expected, as Predetermined widgets impose more limitations and are less flexible for workarounds compared to Open and Controlled widgets.
Data with a score of 1, indicating significant loss in misalignment data, is more common in both Predetermined and Open categories. This often results from the single-widget effect discussed earlier, where users prefer to input a single value when the form provides only one widget (e.g., a combo box or text box) that is conventionally meant to capture a single data item.
We conducted a Kruskal-Wallis H test to determine if there were statistically significant differences in workaround quality across the three widget categories. Although it found a statistically significant difference among the categories ($H(2)=8.37,p=0.015$), the post hoc pairwise comparisons using the Mann-Whitney U test with Bonferroni correction ($p<0.017$) only reported significant differences between Predetermined and Controlled widget categories. Therefore, the integrity of the data is better preserved when employing Controlled widgets in workarounds.
Answer to RQ2: Users often rely on open widgets when employing Fallback methods. Their flexibility allows users to input unstructured or misaligned data, supplement other workaround types, and even repurpose them for unintended applications. Predetermined widgets are often used for Functional Adjustments but are more prone to Data Loss due to their limited flexibility. Controlled widgets, despite their constraints, enable higher-quality data entry through creative user adaptations and complementary use of open widgets.
4.3 Effort and Systematic Use of Workarounds (RQ3)
We now report on the user’s behavior when encountering the same misalignment types multiple times.
Time effort. We assess participant effort using the relative workaround effort (Section 3.0.5), measuring how much longer or shorter it took to save a misaligned record compared to baseline save time. We used Spearman’s test to examine the correlation between relative workaround effort and the order of repeated misalignment types. Our results showed a significant weak negative correlation between the two variables $(r_{s}=-0.377,p=0.000,n=898)$. This indicates that the more often participants encounter a specific type of misalignment, the less time they spend saving subsequent records of that type. We can infer that participants recognize recurring patterns and know how to handle the misalignment the more often they encounter it.
Systematicity. To evaluate whether participants employed the same workarounds systematically when encountering the same misalignment type, we rely on the systematicity score (see Section 3.0.5). Across all participants and apps ($n=45$), 88.1% obtained a perfect score of 1, while the remaining case scored at least 0.87. This indicates that participants consistently used the same workaround type for the same misalignment type.
Quality consistency. To assess the consistency of workaround quality, we calculated the interquartile range (IQR) of workaround quality scores for each participant in each app. Our results showed that 90.5% of the participants had an IQR1absent1$\leq$ 1≤ 1. The maximum $IQR$ observed was 2 for one participant in one app. The mean IQR is 0.74 indicating that participants maintained a consistent quality level when addressing different misalignments.
Survey results. The survey results further supported these findings. 80% of participants agreed or strongly agreed with the statement, “I found another way to enter the records with difficulties (a workaround)”, indicating they employed workarounds when encountering misalignments. Additionally, 93.3% agreed or strongly agreed with, “I used the same workaround for records with similar difficulties”, highlighting a systematic approach, as participants consistently applied the same type of workaround to similar misalignments.
Answer to RQ3: As users encounter the same misalignment type more frequently, they spend less time working around it. This behaviour suggests that users develop systematic strategies and once they successfully apply a workaround type for a specific misalignment, they tend to reuse it for similar misalignments in the future while maintaining a consistent level of quality.
5 Discussion
Our findings reveal distinct trends in how users handle misalignments in software application forms. Users strongly prefer workarounds over abandoning data entry, prioritizing strategies that preserve data integrity. Functional and Data Adjustments were the most frequent and effective workaround types, often resulting in higher data quality compared to Fallback or Data Loss. Users frequently combined multiple workaround types, particularly pairing Functional and Data Adjustments with Fallback methods to document additional details. While Fallback methods were less effective alone, they complemented other strategies by allowing users to capture otherwise unsupported information.
Users demonstrated a systematic approach when encountering repeated issues, becoming more efficient over time. A strong correlation between workaround type and data quality underscores their prioritization of accuracy. Open widgets like text boxes offered flexibility but were more prone to data loss due to users assuming one widget can only capture one piece of data. Predetermined widgets, like checkboxes, encouraged creative Functional Adjustments, while Controlled widgets, such as date pickers, facilitated higher data quality despite their constraints by guiding users toward more disciplined strategies. This challenges the assumption that more flexible widgets inherently result in better data capture, suggesting instead that well-defined constraints guide users toward disciplined, higher-quality strategies for workarounds.
These findings confirm the dual nature of workarounds. One the one hand, they expose design flaws, such as missing features or misaligned workflows. On the other hand, they pose risks to data integrity, especially when Fallback or Data Loss strategies are used. Therefore, it is crucial that software designers proactively understand workarounds to identify unmet user needs and mitigate their risks through timely detection. The results from our survey confirms that detecting workarounds and their reasons can uncover design problems. For instance, participants suggested features like “multi-option selectable for the dropdown” or “more flexibility in attribute choices (topic of interest)”, assuming combo boxes were limited to one piece of data, even though the designer intended the combo box to be editable, allowing users to add new values as needed. Early detection of such workarounds can address these issues. Another comment, “No need for manual input when Excel does the job”, illustrates how designs unaligned with users’ needs can drive them to adopt shadow systems or complete bypass [9], supplementing or replacing the main system with alternatives.
6 Threats to Validity
Construct validity. One threat to validity is whether our workaround quality measurement accurately reflects true quality. While we based it on data retention to reduce subjectivity, it may not fully capture overall quality. Similarly, the systematicity score assumes that repeated use of the same workaround type indicates intentional behaviour, which may be influenced by factors like user background or problem interpretation.
Internal validity. Participants may have attempted more workarounds than usual due to the Hawthorne effect [1]. To mitigate this, we carefully worded the invitation and instructions of the experiment to focus on data entry and software interaction, avoiding any mention of workarounds. Additionally, while tasks and misalignments were randomized, earlier tasks in some sequences may have unintentionally helped participants perform better in later tasks.
External validity. The tasks, application domains, and synthetic data used in our study may not fully represent real-world scenarios or capture the complexity of real-world misalignments, limiting the generalizability of our results. While forms are pervasive in software and provide a suitable focus for this stage of research, the form-based scope may limit the applicability of our findings to other systems where workarounds may differ. The participants’ engineering and computer science background may have influenced their approach to workarounds, affecting the generalizability of our findings. We tried to mitigate this by assigning them non-technical tasks, however it is possible that, e.g., computational thinking [26] might influence the types of workarounds employed. Repeating the experiment with more diverse participants could help address this limitation.
Conclusion validity. We had a small number of participants, which reduced the statistical power of our analysis. While we identified significant associations, the moderate strength of some statistical tests limits the generalizability of our findings to larger datasets and populations. Finally, our interpretations of why participants chose certain workaround types or widgets may involve subjective judgment. To mitigate this, we first interpreted the results and then cross-referenced participants’ responses from the exit survey to corroborate our findings.
7 Conclusion
We investigated how users employ workarounds when software forms do not align with their data or goals. We collected a dataset of 920 user workarounds, and analyzed it for patterns in workaround types, their combinations, and their effects on system design and data quality. We contribute insights into user behaviour when faced with misalignments, a publicly available dataset, and findings to support future research and guide designers in addressing user needs. Future work could explore non-form-based systems, include more diverse participants, and develop automated workaround detection to create more user-centered software.